# Learning to Defer in Content Moderation: The Human-AI Interplay

### v2

@@ C.4 Regret guarantee for OLBACID (Theorem 3)
$\sum_{t=1}^{T}E_{k}(t)=\sum_{t=1}^{T_{k}-1}E_{k}(t)+1\leq n_{k}(T_{k})+1\leq\frac{32\sigma_{\max}^{2}\ln T}{\max(\eta,\gamma)^{2}}+1\leq\frac{33\sigma_{\max}^{2}\ln T}{\max(\eta,\gamma)^{2}}.$
Summing across all types gives $\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\leq\frac{33K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}}$. ∎
C.7 Lipschitz properties of functions $r^{O},r^{R},r$ (Lemma 4.8)
@@
We fix the type $k$ and recall the definition of $r_{k}^{O}$ and $r_{k}^{R}$ in (5). We first show the Lipschitz property for $r_{k}^{O}$. By definition (5), we have
$\left|r_{k}^{O}(h_{1})-r_{k}^{O}(h_{2})\right|\leq v_{k}|h_{1}\Phi(h_{1}/\sigma_{k})-h_{2}\Phi(h_{2}/\sigma_{k})|+v_{k}\sigma_{k}|\varphi(h_{1}/\sigma_{k})-\varphi(h_{2}/\sigma_{k})|.$ (19)
Let $f(x)=x\Phi(x/\sigma_{k}).$ Then $f^{\prime}(x)=\Phi(x/\sigma_{k})+(x/\sigma_{k})\varphi(x/\sigma_{k})=\Phi(x/\sigma_{k})-\varphi^{\prime}(x/\sigma_{k})$ where we use the fact that the derivative of $\varphi(y)$, $\varphi^{\prime}(y)$, is given by $-y\varphi(y)$. By Fact 6, we have $|\varphi^{\prime}(y)|\leq 1/4$ for any $y$. As a result, $|f^{\prime}(x)|\leq|\Phi(x/\sigma_{k})|+|\varphi^{\prime}(x/\sigma_{k})|\leq 1+1/4=5/4$ and thus $f(x)$ is $5/4$-Lipschitz continuous. This shows that $v_{k}|h_{1}\Phi(h_{1}/\sigma_{k})-h_{2}\Phi(h_{2}/\sigma_{k})|\leq\frac{5}{4}v_{k}|h_{2}-h_{1}|$.
In addition, since $|\varphi^{\prime}(y)|\leq 1/4$, we know $\varphi(y)$ is $1/4$-Lipschitz continuous, which shows $v_{k}\sigma_{k}|\varphi(h_{1}/\sigma_{k})-\varphi(h_{2}/\sigma_{k})|\leq v_{k}\sigma_{k}/4|h_{1}-h_{2}|/\sigma_{k}=\frac{1}{4}v_{k}|h_{1}-h_{2}|$. As a result of (19), we have $|r_{k}^{O}(h_{1})-r_{k}^{O}(h_{2})|\leq\frac{5}{4}v_{k}|h_{1}-h_{2}|$ and thus $r_{k}^{O}(\cdot)$ is $\frac{3}{2}v_{k}$-Lipschitz continuous. Since $r_{k}^{R}(h)=r_{k}^{O}(-h)$, we also show the Lipschitz property for $r_{k}^{R}$.
We next show the property for $r_{k}(\cdot)$. Let $h_{1}\leq h_{2}\in[-H,H]$. We need to bound $|r_{k}(h_{1})-r_{k}(h_{2})|$. There are three cases:
  • $h_{1}\leq h_{2}\leq 0$. Then $r_{k}(h_{1})=r_{k}^{O}(h_{1}),r_{k}(h_{2})=r_{k}^{O}(h_{2})$. Since $r_{k}^{O}$ is $(5/4)v_{k}$-Lipschitz, we have $|r_{k}(h_{1})-r_{k}(h_{2})|\leq(3/2)v_{k}|h_{1}-h_{2}|$;
  • $0\leq h_{1}\leq h_{2}$, same as in the last case but now we have $r_{k}(h_{1})=r_{k}^{R}(h_{1}),r_{k}(h_{2})=r_{k}^{R}(h_{2})$;
  • $h_{1}<0<h_{2}$. We have $r_{k}(h_{1})=r_{k}^{O}(h_{1})$ and $r_{k}(h_{2})=r_{k}^{R}(h_{2})$, so $|r_{k}(h_{1})-r_{k}(h_{2})|=|r_{k}^{O}(h_{1})-r_{k}^{R}(h_{2})|$. In addition,
    $|r_{k}^{O}(h_{1})-r_{k}^{R}(h_{2})|$ $=v_{k}\left|h_{1}\Phi(h_{1}/\sigma_{k})+\sigma_{k}\varphi(h_{1}/\sigma_{k})+h_{2}\Phi(-h_{2}/\sigma_{k})-\sigma_{k}\varphi(-h_{2}/\sigma_{k})\right|$
    $\leq v_{k}\left(\sigma_{k}|\varphi(h_{1}/\sigma_{k})-\varphi(-h_{2}/\sigma_{k})|+|h_{1}\Phi(h_{1}/\sigma_{k})+h_{2}\Phi(-h_{2}/\sigma_{k})|\right).$
    We next bound the two terms in the parenthesis. For the first term, since $\varphi(y)$ is $1/4-$Lipschitz, we have $\sigma_{k}|\varphi(h_{1}/\sigma_{k})-\varphi(-h_{2}/\sigma_{k})|\leq\frac{1}{4}|h_{1}+h_{2}|\leq(h_{2}-h_{1})/4$ since $h_{1}<0<h_{2}$. For the second term, recall that $f(x)=x\Phi(x/\sigma_{k})$ and $f(x)$ is $5/4$-Lipschitz. We have $|h_{1}\Phi(h_{1}/\sigma_{k})+h_{2}\Phi(-h_{2}/\sigma_{k})|=f(h_{1})-f(-h_{2})\leq(5/4)|h_{1}+h_{2}|\leq(5/4)(h_{2}-h_{1})$. Summing the two terms give $|r_{k}(h_{1})-r_{k}(h_{2})|=|r_{k}^{O}(h_{1})-r_{k}^{R}(h_{2})|\leq(3/2)v_{k}(h_{2}-h_{1})=(3/2)v_{k}|h_{1}-h_{2}|.$
Summing the three cases shows that $r_{k}(\cdot)$ is $\frac{3}{2}v_{k}$-Lipschitz continuous, completing the proof. ∎
C.8 Bounding per-period classification loss (Lemma 4.9)
Proof of Lemma 4.9).
@@
    $\frac{Z_{k}(t)}{\ell_{k}}\leq|r_{k}^{R}-r_{k}^{O}|$ $=|r_{k}^{R}(h_{k})-r_{k}^{O}(h_{k})|$
    $\leq|r_{k}^{R}(0)-r_{k}^{O}(0)|+|r_{k}^{R}(h_{k})-r_{k}^{R}(0)|+|r_{k}^{O}(h_{k})-r_{k}^{O}(0)|$
    $\leq|r_{k}^{R}(h_{k})-r_{k}^{R}(0)|+|r_{k}^{O}(h_{k})-r_{k}^{O}(0)|\leq 3v_{k}|h_{k}|\leq 3v_{k}\gamma,$ (20)
    where the second inequality uses $r_{k}^{R}(0)=r_{k}^{O}(0)$; the third inequality uses the Lipschitz property in Lemma 4.8; and the last inequality uses $h_{k}\in[-\gamma,0]$.
    Alternatively, if $\hat{h}_{k}(t)\leq 0$, we set $Y_{k}(t)=1$. By definition of the confidence interval (11), we have $\bar{h}_{k}(t)\leq\hat{h}_{k}(t)+(\hat{h}_{k}(t)-\underline{h}_{k}(t))\leq\gamma$ because $\underline{h}_{k}(t)\geq-\gamma\geq-H$ (recall that we assume $H\geq 1\geq\gamma$). Since the confidence bound is valid conditioning on $\mathcal{E}_{k,t}$, we have $h_{k}\in[-\gamma,\gamma]$. As a result, $Z_{k}(t)/\ell_{k}\leq|r_{k}^{R}-r_{k}^{O}|\leq 3v_{k}|h_{k}|\leq 3v_{k}\gamma$ as in (20).
  • $\bar{h}_{k}(t)\leq\gamma$; This case is symmetric to the above case. Following the same analysis gives $Z_{k}(t)\leq 3v_{k}\gamma\ell_{k}$ as well.
@@
C.9 Bounding idiosyncrasy loss by Lagrangian via drift analysis (Lemma 4.10)
@@
C.10 Bounding the Regret in Admission for OLBACID (Lemma 4.11)
@@
We have $\sum_{t=1}^{T}\bar{A}_{k}(t)\min\left(r_{\max},3v_{k}\sigma_{k}\sqrt{\frac{8\ln t}{n_{k}(t)}}\right)\leq 4\beta r^{2}_{\max}\ell_{\max}+6v_{k}\sigma_{k}\sqrt{8T\ln T}$.
@@
Recall that $A_{k}(t)$ captures whether the algorithm admits a type-$k$ post into the review queue and $E_{k}(t)$ captures whether it admits a type-$k$ post into the label-driven queue. By Line 2 of Algorithm 2, we know that $A_{k}(t)=\bar{A}_{k}(t)(1-E_{k}(t))$ and thus $\bar{A}_{k}(t)\leq A_{k}(t)+E_{k}(t)$. We thus aim to upper bound $\sum_{t=1}^{T}(A_{k}(t)+E_{k}(t))\min\left(r_{\max},3v_{k}\sigma_{k}\sqrt{\frac{8\ln t}{n_{k}(t)}}\right)$. Denoting the sequence of periods with $A_{k}(t)+E_{k}(t)=1$ as $t_{1},\ldots,t_{Y}$ where $t_{Y}=\sum_{t=1}^{T}A_{k}(t)+E_{k}(t)$ and setting $n^{\prime}(y)=n_{k}(t_{y})$, we then need to upper bound $\sum_{y=1}^{Y}\min\left(r_{\max},3v_{k}\sigma_{k}\sqrt{\frac{8\ln T}{n^{\prime}(y)}}\right).$
@@
$\sum_{t=1}^{T}\bar{A}_{k}(t)\min\left(r_{\max},3v_{k}\sigma_{k}\sqrt{\frac{8\ln t}{n_{k}(t)}}\right)$ $\leq\sum_{y=1}^{Y}\min\left(r_{\max},3v_{k}\sigma_{k}\sqrt{\frac{8\ln T}{n^{\prime}(y)}}\right)$
$\leq\sum_{y=1}^{Y}\min\left(r_{\max},3v_{k}\sigma_{k}\sqrt{\frac{8\ln T}{\max(0,y-4\beta r_{\max}\ell_{\max})}}\right)$
$\overset{u=y-4\beta r_{\max}\ell_{\max}}{\leq}4\beta r^{2}_{\max}\ell_{\max}+\sum_{u=1}^{T}3v_{k}\sigma_{k}\sqrt{\frac{8\ln T}{u}}\leq 4\beta r^{2}_{\max}\ell_{\max}+6v_{k}\sigma_{k}\sqrt{8T\ln T}.$
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}(A_{k}^{\textsc{BACID}}(t)-\bar{A}_{k}(t))(\beta r_{k}\ell_{k}-Q_{k}(t))\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\right]+4\beta Kr_{\max}\ell_{\max}+4K$ (21)
where the first inequality uses $Q_{k}(t)\leq t$ and the second inequality uses Lemma 4.5 that $\mathbb{P}\{\mathcal{E}_{k,t}\}\geq 1-2t^{-3}$ and that $\sum_{t=1}^{T}t^{-2}\leq 2$.
We fix a type $k$ and upper bound $\sum_{t=1}^{T}(A_{k}^{\textsc{BACID}}(t)-\bar{A}_{k}(t))(\beta r_{k}\ell_{k}-Q_{k}(t))\mathbbm{1}\left(\mathcal{E}_{k,t}\right)$. Fixing a period $t$ and assuming $\mathcal{E}_{k,t}$ holds, we have $h_{k}\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]$ and thus $\bar{r}_{k}(t)=\max_{h\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]}r_{k}(h)\geq r_{k}(h_{k})=r_{k}$. As a result,
$\bar{r}_{k}(t)-r_{k}=\max_{h\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]}r_{k}(h)-r_{k}(h_{k})\leq\frac{3}{2}v_{k}\left(\bar{h}_{k}(t)-\underline{h}_{k}(t)\right)\leq 3v_{k}\sigma_{k}\sqrt{\frac{8\ln t}{n_{k}(t)}}$ (22)
where the first inequality uses the Lipschitz property of $r_{k}(\cdot)$ in Lemma 4.8 and the fact that $h_{k}\in[\underline{h}_{k}(t),\bar{h}_{k}(t)]$; the second inequality uses (11). In addition, if $\Lambda_{k}(t)=0$, we have $\bar{A}_{k}(t)=A_{k}^{\textsc{BACID}}(t)=0$; otherwise,
@@
$\leq\beta(\bar{r}_{k}(t)-r_{k})\ell_{k}\leq\beta\min\left(r_{\max},3v_{k}\sigma_{k}\sqrt{\frac{8\ln t}{n_{k}(t)}}\right)\ell_{k}$
where the last inequality is because $\bar{r}_{k}(t)\leq r_{\max}$ by $\underline{h}_{k}(t),\bar{h}_{k}(t)\in[-H,H]$ and (22). Therefore,
$\sum_{t=1}^{T}(A_{k}^{\textsc{BACID}}(t)-\bar{A}_{k}(t))(\beta r_{k}\ell_{k}-Q_{k}(t))\mathbbm{1}\left(\mathcal{E}_{k,t}\right)$ $\leq\beta\ell_{k}\sum_{t=1}^{T}\bar{A}_{k}(t)\min\left(r_{\max},3v_{k}\sigma_{k}\sqrt{\frac{8\ln t}{n_{k}(t)}}\right)$
@@
$\textsc{RegA}(T)$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}(A_{k}^{\textsc{BACID}}(t)-\bar{A}_{k}(t))(\beta r_{k}\ell_{k}-Q_{k}(t))\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\right]+4\beta Kr_{\max}\ell_{\max}+4K$
$\leq 4K\beta^{2}r^{2}_{\max}\ell^{2}_{\max}+6K\beta v_{\max}\sigma_{\max}\ell_{\max}\sqrt{8T\ln T}+4\beta Kr_{\max}\ell_{\max}+4K$
$\leq 12K\beta^{2}r^{2}_{\max}\ell^{2}_{\max}+6K\beta v_{\max}\sigma_{\max}\ell_{\max}\sqrt{8T\ln T}$
$\leq 12K\beta^{2}r^{2}_{\max}\ell^{2}_{\max}+6K\beta r_{\max}\ell_{\max}\sqrt{8T\ln T}$
where we use the assumption that $K\leq\ell_{\max}$, $\beta\geq 1/\ell_{\max}$ and $r_{\max}=v_{\max}(H+\sigma_{\max})$. ∎
C.11 Bounding the Regret in Scheduling for OLBACID (Lemma 4.12)
@@
$\leq 2\beta r_{\max}\ell_{\max}\mathbb{E}\left[\sum_{t=1}^{T}Q^{\mathcal{\textsc{LD}}}(t)\right]\leq\frac{66\beta r_{\max}\ell_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}$

### v3

@@ C.3 Regret guarantee for OLBACID (Theorem 3)
$\sum_{t=1}^{T}E_{k}(t)\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\leq\sum_{t=1}^{T_{k}-1}E_{k}(t)+1\leq n_{k}(T_{k})+1\leq\frac{32\sigma_{\max}^{2}\ln T}{\max(\eta,\gamma)^{2}}+1\leq\frac{33\sigma_{\max}^{2}\ln T}{\max(\eta,\gamma)^{2}}.$
Summing across all types gives $\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\leq\frac{33K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}}$. Therefore,
$\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)(\mathbbm{1}\left(\mathcal{E}_{k,t}+\mathcal{E}_{k,t}^{c}\right))\right]\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}E_{k}(t)\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\right]+5\leq\frac{33K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}}+5\leq\frac{38K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}}.$
C.6 Bounding per-period classification loss (Lemma 4.8)
@@
    $\frac{Z_{k}(t)}{\ell_{k}}\leq|r_{k}^{R}-r_{k}^{O}|=\left|\mathbb{E}_{k}\left[-c_{j}^{-}\right]-\mathbb{E}_{k}\left[c_{j}^{+}\right]\right|=|-h_{k}|\leq\gamma,$ (22)
    where the last inequality uses $h_{k}\in[-\gamma,0]$.
    Alternatively, if $\hat{h}_{k}(t)\leq 0$, we set $Y_{k}(t)=1$. By definition of the confidence interval (11), we have $\bar{h}_{k}(t)\leq\hat{h}_{k}(t)+(\hat{h}_{k}(t)-\underline{h}_{k}(t))\leq\gamma$ because $\underline{h}_{k}(t)\geq-\gamma\geq-H$ (recall that we assume $H\geq 1\geq\gamma$). Since the confidence bound is valid conditioning on $\mathcal{E}_{k,t}$, we have $h_{k}\in[-\gamma,\gamma]$. As a result, $Z_{k}(t)/\ell_{k}\leq|r_{k}^{R}-r_{k}^{O}|=|-h_{k}|\leq\gamma$ as in (22).
  • $\bar{h}_{k}(t)\leq\gamma$; This case is symmetric to the above case. Following the same analysis gives $Z_{k}(t)\leq\gamma\ell_{k}$ as well.
@@
C.7 Bounding idiosyncrasy loss by Lagrangian via drift analysis (Lemma 4.9)
@@
C.8 Bounding the Regret in Admission for OLBACID (Lemma 4.10)
@@
We have $\sum_{t=1}^{T}\bar{A}_{k}(t)\min\left(r_{\max},4\sigma_{\max}\sqrt{\frac{\ln t}{n_{k}(t)}}\right)\leq 4\beta r^{2}_{\max}\ell_{\max}+8\sigma_{\max}\sqrt{T\ln T}$.
@@
Recall that $A_{k}(t)$ captures whether the algorithm admits a type-$k$ post into the review queue and $E_{k}(t)$ captures whether it admits a type-$k$ post into the label-driven queue. By Line 2 of Algorithm 2, we know that $A_{k}(t)=\bar{A}_{k}(t)(1-E_{k}(t))$ and thus $\bar{A}_{k}(t)\leq A_{k}(t)+E_{k}(t)$. We thus aim to upper bound $\sum_{t=1}^{T}(A_{k}(t)+E_{k}(t))\min\left(r_{\max},4\sigma_{\max}\sqrt{\frac{\ln t}{n_{k}(t)}}\right)$. Denoting the sequence of periods with $A_{k}(t)+E_{k}(t)=1$ as $t_{1},\ldots,t_{Y}$ where $t_{Y}=\sum_{t=1}^{T}A_{k}(t)+E_{k}(t)$ and setting $n^{\prime}(y)=n_{k}(t_{y})$, we then need to upper bound $\sum_{y=1}^{Y}\min\left(r_{\max},4\sigma_{\max}\sqrt{\frac{\ln T}{n^{\prime}(y)}}\right).$
@@
$\sum_{t=1}^{T}\bar{A}_{k}(t)\min\left(r_{\max},4\sigma_{\max}\sqrt{\frac{\ln t}{n_{k}(t)}}\right)$ $\leq\sum_{y=1}^{Y}\min\left(r_{\max},4\sigma_{\max}\sqrt{\frac{\ln T}{n^{\prime}(y)}}\right)$
$\leq\sum_{y=1}^{Y}\min\left(r_{\max},4\sigma_{\max}\sqrt{\frac{\ln T}{\max(0,y-4\beta r_{\max}\ell_{\max})}}\right)$
$\overset{u=y-4\beta r_{\max}\ell_{\max}}{\leq}4\beta r^{2}_{\max}\ell_{\max}+\sum_{u=1}^{T}4\sigma_{\max}\sqrt{\frac{\ln T}{u}}\leq 4\beta r^{2}_{\max}\ell_{\max}+8\sigma_{\max}\sqrt{T\ln T}.$
@@
$\leq\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}(A_{k}^{\textsc{BACID}}(t)-\bar{A}_{k}(t))(\beta r_{k}\ell_{k}-Q_{k}(t))\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\right]+8\beta Kr_{\max}\ell_{\max}+8K$ (23)
where the first inequality uses $Q_{k}(t)\leq t$ and the second inequality uses Lemma 4.5 that $\mathbb{P}\{\mathcal{E}_{k,t}\}\geq 1-4t^{-3}$ and that $\sum_{t=1}^{T}t^{-2}\leq 2$.
We fix a type $k$ and upper bound $\sum_{t=1}^{T}(A_{k}^{\textsc{BACID}}(t)-\bar{A}_{k}(t))(\beta r_{k}\ell_{k}-Q_{k}(t))\mathbbm{1}\left(\mathcal{E}_{k,t}\right)$. Fixing a period $t$ and assuming $\mathcal{E}_{k,t}$ holds, we have $r_{k}\leq\bar{r}_{k}(t)$ and $\bar{r}_{k}(t)-r_{k}=4\sigma_{\max}\sqrt{\frac{\ln t}{n_{k}(t)}}$ by the definition of $\bar{r}_{k}(t)$ in (12). In addition, if $\Lambda_{k}(t)=0$, we have $\bar{A}_{k}(t)=A_{k}^{\textsc{BACID}}(t)=0$; otherwise,
@@
$\leq\beta(\bar{r}_{k}(t)-r_{k})\ell_{k}\leq\beta\min\left(r_{\max},4\sigma_{\max}\sqrt{\frac{\ln t}{n_{k}(t)}}\right)\ell_{k}$
where the last inequality uses the definition of $\bar{r}_{k}(t)$ in (12). Therefore,
$\sum_{t=1}^{T}(A_{k}^{\textsc{BACID}}(t)-\bar{A}_{k}(t))(\beta r_{k}\ell_{k}-Q_{k}(t))\mathbbm{1}\left(\mathcal{E}_{k,t}\right)$ $\leq\beta\ell_{k}\sum_{t=1}^{T}\bar{A}_{k}(t)\min\left(r_{\max},4\sigma_{\max}\sqrt{\frac{\ln t}{n_{k}(t)}}\right)$
@@
$\textsc{RegA}(T)$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}\sum_{k\in\mathcal{K}}(A_{k}^{\textsc{BACID}}(t)-\bar{A}_{k}(t))(\beta r_{k}\ell_{k}-Q_{k}(t))\mathbbm{1}\left(\mathcal{E}_{k,t}\right)\right]+8\beta Kr_{\max}\ell_{\max}+8K$
$\leq 4K\beta^{2}r^{2}_{\max}\ell^{2}_{\max}+8K\beta\sigma_{\max}\ell_{\max}\sqrt{T\ln T}+8\beta Kr_{\max}\ell_{\max}+8K$
$\leq 20K\beta^{2}r^{2}_{\max}\ell^{2}_{\max}+8K\beta\sigma_{\max}\ell_{\max}\sqrt{T\ln T},$
where we use the assumption that $K\leq\ell_{\max}$, $\beta\geq 1/\ell_{\max}$ and $r_{\max}\geq 1$. ∎
C.9 Bounding the Regret in Scheduling for OLBACID (Lemma 4.11)
@@
$\leq 2\beta r_{\max}\ell_{\max}\mathbb{E}\left[\sum_{t=1}^{T}Q^{\mathcal{\textsc{LD}}}(t)\right]\leq\frac{76\beta r_{\max}\ell_{\max}K\sigma^{2}_{\max}\ln T}{\max(\eta,\gamma)^{2}\hat{\mu}_{\min}}$